OpenStudy (anonymous):

Solve the following system of equations. Round your answers to the nearest tenth. Your answer will be in the form: (M,N). Find M + N, rounded to the nearest tenth. 5x+2y=21 -2x+6y=-34

OpenStudy (anonymous):

A) 5x + 2y = 21 B) -2x + 6y = -34 2xA = ? 5xB = ?

OpenStudy (anonymous):

say...

OpenStudy (anonymous):

wait 2xA.... and 5xB?

OpenStudy (anonymous):

yes I meant this: \[2\times(5x+2y) = 2\times(21)\] and this: \[5\times(-2x+6y)= 5\times (-34)\]

OpenStudy (anonymous):

okay let me try it

OpenStudy (anonymous):

10x+4y=42 -10x+30y=-170 like that?

OpenStudy (anonymous):

yes! :) and now find this: A + B = ?

OpenStudy (anonymous):

D: okay

OpenStudy (anonymous):

how am i going to erase the x variable when it is already a negative?!?! DDD:::

OpenStudy (anonymous):

would it be 0x+34y=-128?

OpenStudy (anonymous):

in this way: A + B = (10x-10x) + 34y = -128 -> 34y = -128

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

so find y...

OpenStudy (anonymous):

y=-3.76470588

hartnn (hartnn):

just to give an alternate way: multiply A) by 3 and subtract the equations..

OpenStudy (anonymous):

yes,@hartnn said...and you found y,so you can find x .. ;)

OpenStudy (anonymous):

okay!!

hartnn (hartnn):

so could you find x ?

OpenStudy (anonymous):

5x+2(-3.765)=21 =5x+(-7.53)+7.52941176=21+7.53 =5x=28.53 =x=5.706

OpenStudy (anonymous):

and the last part for (M,N) M+N So (5.706, -3.765) and M+N = 5.706+(-3.765)= 1.941 ~ 1.9

OpenStudy (anonymous):

it is correct?

hartnn (hartnn):

yes :D

OpenStudy (anonymous):

:DDD yay!! thank you both

OpenStudy (anonymous):

you're welcome madam...:)